OpenStudy (anonymous):

when simplify what is 4 over 6-6?

jimthompson5910 (jim_thompson5910):

is 6-6 in the denominator or is it just one single 6?

myininaya (myininaya):

\[\frac{4}{6}-6?\]

OpenStudy (anonymous):

its the demominator

myininaya (myininaya):

\[\frac{4}{6-6} does \not exist \]

OpenStudy (anonymous):

yeah how myininaya wrote it

myininaya (myininaya):

whenever you have 0 on the bottom the number does not exist

jimthompson5910 (jim_thompson5910):

so then it's (4/6) - 6 = (2/3) - 6 = (2/3)-(18/3) = (2-18)/3 = -16/3 so the answer is -16/3

myininaya (myininaya):

if you were talking about the first way i wrote the answer is -16/3 the second way i wrote it the number does not exist